# Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework

### v2

@@ 5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
After $T$ optimization steps, the sequence of best-performing outcomes $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$, where $j_{t}^{\ast}:=\operatorname{argmax}_{j}s_{t}^{j}$, is reported as the algorithm’s result.
Note that T-BoN BO takes $N\times G$ generation of prompt/outcome candidates before each evaluation. This careful selection of a prompt/outcome candidate to evaluate is the key to evaluation efficiency in T-BoN BO: it deliberately trades inexpensive generation compute for fewer evaluations.
Input: System $\Phi$; $\{\Pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$; $J$ trajectories Output: Optimized triples $\{(\Pi^{j}_{T},c^{j}_{T},s^{j}_{T})\}_{j=1}^{J}$ 1 for $j=1$ to $J$ do 2 $c_{0}^{j}\leftarrow\Phi(\Pi_{0}^{j})$; 3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{0}))$; 4 $H_{0}\leftarrow\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$, $R\leftarrow\emptyset$; 5  6 end for 7 8for $t=1$ to $T$ do 9 for $j=1$ to $J$ do 10    $\Pi^{j}_{t,0}\leftarrow\Pi^{j}_{t-1}$; 11    for $g=1$ to $G$ do 12       for $i=1$ to $N$ do 13          $\delta_{t,g}^{j,(i)}\leftarrow\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R,M_{\text{critic}})$ ; 14          $\Pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\Pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)})$ ; 15          $c_{t,g}^{j,(i)}\leftarrow\Phi(\Pi_{t,g}^{j,(i)})$ ; 16           17         end for 18       $(\Pi_{t,g}^{j},c_{t,g}^{j})\leftarrow\textsc{Best-of-N}(\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\},R_{t-1},M_{\text{critic}})$; 19        20      end for 21    $c_{t}^{j}\leftarrow c^{j}_{t,G}$, $\Pi_{t}^{j}\leftarrow\Pi_{t,G}^{j}$; 22    $s_{t}^{j}\leftarrow\text{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}))$; 23    $H_{t}\leftarrow H_{t-1}\cup\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}$; 24    if $s_{t}^{j}<s_{t-1}^{j}$ then 25       $\Pi_{t}^{j}\leftarrow\Pi_{t-1}^{j}$, $c_{t}^{j}\leftarrow c_{t-1}^{j}$, $s_{t}^{j}\leftarrow s_{t-1}^{j}$; 26        27      end if 28     29   end for 30 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$; 31  32 end for return $\{c_{t}^{j_{t}^{*}}\}_{t=1}^{T}$, where $j_{t}^{*}:=\operatorname{argmax}_{j}s_{t}^{j}$ Algorithm 1 T-BoN BO
6 Case Study: Ad Optimization in Digital Marketing
[…]
Figure 5: An ad generation AI system $\Phi$ and prompt $\Pi=\langle\pi\rangle$.
[…]
Figure 6: A self-improving AI implemented on the ad generation AI system in Figure 5.
In this section, we present a case study on how ad platforms can use T-BoN BO for ad optimization and automate the ad generation, analysis, and evaluation cycle. Given an a priori fine-tuned image generation AI model that is aligned with its client’s brand, the objective is to efficiently refine the prompts for ad generation to better align with the users’ preference distribution.
6.1 Problem definition
Formally, define the brand fine-tuned ad-generation AI system as $\Phi$, whose behavior is controlled by a prompt $\Pi$ (Figure 5). For a given prompt $\Pi$, we write $\Phi(\Pi)$ for the corresponding ad creative shown on the ad platform. As a system, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ (potential ad viewers on the platform) to a distribution over the output space $\mathcal{Y}$ (observable ad responses). For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued output of the system at input $x$. In a digital marketing setting, $x$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$. When a user $x\in\mathcal{X}$ is shown the ad creative $\Phi(\Pi)$, we observe an outcome $y\in\mathcal{Y}$, which is a realization of $y(\Phi(\Pi),x)$.
As discussed earlier, the distribution of exposed users depends on the ad $\Phi(\Pi)$ and therefore on the prompt $\Pi$. We denote by $\mathcal{D}_{\mathcal{X}}(\Pi)$ the unknown distribution of users $x$ who are shown the ad generated from $\Pi$ when the ad $\Phi(\Pi)$ is run on the platform.
To evaluate system performance, a marketer pre-defines a scoring rule $r:\mathcal{Y}\to[0,1]$ that maps observed user responses (clicks, conversions, spend) to a normalized effectiveness score. The ad optimization objective of the agency for a given client and campaign can then be written as
$J(\Pi)\;:=\;\mathbb{E}\big[\,r\big(y(\Phi(\Pi),x)\big)\big],$ (23)
where the expectation is taken over $x\sim\mathcal{D}_{\mathcal{X}}(\Pi)$ and any additional randomness in $y(\Phi(\Pi),x)$.
The self–improving ad system built by the agency can therefore be viewed as an iterative algorithm that searches over prompts $\pi$ in order to maximize $J(\pi)$ while minimizing the number of costly ad evaluations on the platform (Figure 6). At iteration $t$, a candidate prompt $\pi_{t}$ is selected, the corresponding ad creative $\Phi(\pi_{t})$ is generated, and the campaign is (conceptually) exposed to a batch of users $\{x_{t}^{(l)}\}_{l=1}^{L}\sim\mathcal{D}_{\mathcal{X}}(\pi_{t})$. The empirical evaluation
$s_{t}\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\big(y(\Phi(\pi_{t}),x_{t}^{(l)})\big)$ (24)
serves as a noisy estimate of $J(\pi_{t})$ and is fed back into the optimization algorithm (T-BoN BO) to update the history $H_{t}$ and the meta–reflection $R_{t}$.
6.2 Application of T-BoN BO
We now describe how the general T-BoN BO framework in Section 5 is instantiated for ad agency’s ad optimization problem. Specifically, we focus on the parallel version of T-BoN BO (Algorithm 1), where $J$ optimization trajectories start from different initial prompts and run simultaneously.
Initialization.
The decision-maker/platform starts from initial prompts $\{\pi_{0}^{j}\}_{j=1}^{J}$, i.e., prompts already in use or high-quality baselines provided by humans. (e.g., “A simple, photorealistic image of a plant-based burger patty. [^5] Note that in the experiments in Section 7, we use Worst5-of-N prompts as initial starting points, which are higher-quality, much more detailed initial prompts; See Appendix A.) For each initial prompt $\pi_{0}^{j}$, the system generates an ad $\Phi(\pi_{0}^{j})$ and obtains an initial score $s_{0}^{j}$ by deploying the ad and observing outcomes. These tuples $\{(\pi_{0}^{j},\Phi(\pi_{0}^{j}),s_{0}^{j})\}_{j=1}^{J}$ form the initial history $H_{0}$ for the parallel T-BoN BO procedure.
Best-of-NN textual-gradient steps.
Within each optimization iteration $t$, T-BoN BO refines each trajectory’s prompt $\pi_{t-1}^{j}$ through $G$ Best-of-N textual-gradient steps (Figure 4). For a given trajectory $j$ and inner step $g$, the critic model $M_{\text{critic}}$ (a large language model such as Gemini 2.5 Flash) receives as context: (i) the current prompt $\pi_{t,g-1}^{j}$, (ii) the current meta-reflection $R_{t-1}$, and (iii) a short description of the brand and campaign scenario. Conditioned on this context, it proposes $N$ small natural-language edits
$\delta_{t,g}^{j,(1)},\ldots,\delta_{t,g}^{j,(N)}\;\sim\;\nabla_{\text{text}}(\pi_{t,g-1}^{j};R_{t-1},M_{\text{critic}}),$ (25)
which form the textual gradients. Each edit is then applied to the current prompt to obtain $N$ candidate prompts
$\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}\big(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)};M_{\text{critic}}\big),\qquad i=1,\dots,N.$ (26)
These textual gradients correspond to small, directed moves in language space that, according to the critic’s judgment, are likely to increase ad effectiveness. For example, conditioned on the reflection $R_{t-1}$ (e.g., “social occasions and visible grill marks tend to perform better than plain pack shots”), the critic $M_{\text{critic}}$ may propose the following textual-gradient candidates:
  1. 1.
  2. 2.
  3. 3.
Each of these textual gradients $\delta^{(i)}$ is then applied to the original prompt to produce candidate prompts. For example, applying the first gradient yields the prompt:
Applying the second gradient yields the prompt:
To select the next prompt that will go forward to the next gradient step among $N$ candidate prompts $\{\pi_{t,g}^{j,(i)}\}_{i=1}^{N}$, T-BoN BO applies the Best-of-$N$ choice rule over the $N$ corresponding candidate ads $\{c_{t,g}^{j,(i)}\}_{i=1}^{N}$. Following liu2025pairjudgermperformbestofn, we implement Best-of-$N$ via a pairwise tournament: in each match, the critic model $M_{\text{critic}}$ compares two creatives side by side and predicts which will perform better, given the campaign goals and the current meta-reflection $R_{t-1}$ (see Figure 11 in Appendix A). After multiple randomized pairwise comparisons, the winner $c_{t,g}^{j}$ and its corresponding prompt $\pi_{t,g}^{j}$ are carried forward to the next gradient step. Repeating this expansion-and-selection process for $g=1,\dots,G$ yields the refined prompt $\pi_{t}^{j}$ and ad $\Phi(\pi_{t}^{j})$ for trajectory $j$ at iteration $t$.
Evaluation as the costly step.
Once all $G$ gradient steps are completed, T-BoN BO performs a single expensive evaluation for each trajectory by deploying the ad $\Phi(\pi_{t}^{j})$ and computing the score $s_{t}^{j}$. In the real ad-agency use case, this evaluation step would correspond to running the creative in an A/B test, multi-armed bandit experiment, or survey-based study, and mapping the resulting performance metrics to a scalar score via the scoring rule $r$. The new triple $(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})$ is appended to the history $H_{t}$, and a simple acceptance rule is applied: if $s_{t}^{j}$ fails to improve upon $s_{t-1}^{j}$, the trajectory rolls back to $(\pi_{t-1}^{j},\Phi(\pi_{t-1}^{j}),s_{t-1}^{j})$.
Meta-reflection shared across trajectories.
After all $J$ trajectories have been evaluated at iteration $t$, the algorithm updates the global meta-reflection $R_{t}=\textsc{MetaReflect}(H_{t})$, as in Section 5. Concretely, $M_{\text{critic}}$ is prompted to review a subset of the best- and worst-performing ads and prompts in $H_{t}$ and to summarize the key patterns and rules that distinguish successful creatives from unsuccessful ones (see Figure 12 in Appendix A). For example, the reflection may emphasize that scenes with social eating and visible grill marks perform better than isolated product shots, or that emphasizing “satisfying” and “juicy” yields higher effectiveness scores. This reflection is then injected into the context of subsequent textual-gradient generation and Best-of-$N$ comparisons, allowing all trajectories to share accumulated knowledge even though they explore different regions of the prompt space.
Reported outcome.
After $T$ iterations, the system returns the sequence of best-performing ads $\{c_{t}^{j_{t}^{\ast}}\}_{t=1}^{T}$, where $j_{t}^{\ast}=\arg\max_{j}s_{t}^{j}$ is the trajectory that achieves the highest score at iteration $t$.
7 Experiments
We validate the empirical effectiveness of the proposed T-BoN BO framework for a real-world digital marketing problem described in 6. Specifically, we examine if and how much T-BoN BO improves outcomes under a fixed evaluation budget compared to state-of-the-art baselines.
7.1 Evaluation Module using Persona data and LLM-based Simulation
To evaluate how well an ad performs on the target population and improve prompts, we need a target population and a way to measure the preferences of that population for a given ad. There are many ways to do this, such as field experiments, consumer surveys, or simulated responses from a pre-defined target population. In our experiments, we use LLM-generated preferences over the Twin-2k-500 persona distribution provided by toubia2025twin. An important advantage of using this persona distribution is that it allows us to test the performance of different algorithms fast without the need for large-scale field experiments or expensive surveys. [^6] Note that naive field deployment cannot be used to test the relative performance of algorithms since an algorithm’s measured advantage would be a mix of its own effect and distributional effects induced by the ad platform’s ad targeting systems.
This dataset includes the data of 2,058 U.S. adults answering 500 survey questions spanning demographics, personality, cognitive ability, economic preferences, classic heuristics-and-biases tasks, and a pricing study. These survey questions of each persona are passed to a multi-modal LLM (Gemini 2.5 Flash) as context so that the LLM can simulate the ad effectiveness for the persona. [^7] While LLM-based persona simulation is often biased and does not reflect real-world human behavior (li2025llm; peng2025mega), it still induces a valid preference distribution for which, as a benchmark, we would like to see how well a method can optimize. We divided the Twin-2k-500 personas into 80% (1647 personas) for the training set and 20% (411 personas) for the test set. Every time we evaluate an ad during the algorithm (i.e., training stage), we randomly sample 200 personas from the training set and simulate the ad’s effectiveness for them. Every time we want to test an ad, such as a final ad output from the algorithm, we simulate its effectiveness for all the test set personas.
7.2 Experiment Setup
@@
We consider eight synthetic ad campaign scenarios that cover a diverse range of products across distinct categories, defined by the creative brief that outlines the strategic and creative direction for each scenario. (For the creative brief for each scenario, see Appendix B.)
@@
  • Scenario 6: “MindGarden," a subscription-based meditation and mindfulness app.
@@
For each scenario, from its creative brief, we generated 64 high-quality initial prompts that were used as input to the state-of-the-art image generation AI model named Imagen 4  [^8] See Vertex AI’s description (https://console.cloud.google.com/vertex-ai/publishers/google/model-garden/imagen-4.0-generate-preview-06-06) for details. to generate the initial 64 high-quality ads. (For prompts, see Figure 13 and 14 in Appendix A.) To maintain a rigorous baseline, we made the 64 initial prompts detailed enough to yield strong ads. This ensures that the algorithm’s subsequent improvements are not explained as trivial quality gains from having initial under-specified prompts. That is, the measured improvements are attributed to the algorithm’s ability to align ads with the underlying persona distribution better.
When simulating an ad’s effectiveness for a persona, we ask the LLM (with the prompt in Figure 7) what the ad’s “effectiveness” [^9] The vagueness of the term “effectiveness” constitutes an evaluation distribution induced by LLM-persona based on what LLM ‘thinks’ about the term, which we target to align the algorithms with. That is, as long as we consistently query the same LLM to predict “effectiveness” using the same prompt, such vagueness is justified. for the persona will be, with the score of [1, 2, 3, 4, 5]. Here, the score of one is the least effective, and the score of five is the most effective. As the LLM’s internal assignment of probability for the scores is available as ‘logprobs’ [^10] See Vertex AI’s logprobs description (https://cloud.google.com/vertex-ai/generative-ai/docs/model-reference/inference#logprobs) for details., we use them to calculate the mean score, and consider it as the evaluation score of the ad for the persona (which is a fraction number like 3.41), instead of considering the actual response score from the LLM (which is an integer number among 1-5). We then take the average of the evaluation scores for the personas to get the final evaluation score of the ad.
Persona prompt for simulating ad effectiveness. PERSONA DATA: $\{$
Which part of the United States do you currently live in? [Answer]
What is the highest level of schooling or degree that you have completed? [Answer]
$\ldots$ (Many other survey questions and answers) $\ldots$
@@
$\ldots$ (Many other survey questions and answers) $\ldots$
Effective Score Scale Definition:
No explanation. Just the score.
Figure 7: Persona simulation prompt for the digital marketing problem.
Among the initial ads, we identified the best initial ad, which we call Best-of-64, and the worst five initial ads, which we call Worst5-of-64, by running a best-arm identification bandit algorithm (russo2016simple) by alternating the best-arm identification objective and worst 5-arm identification objective for evaluations of 5,000 sequential random samples from the training persona set.

### v3

@@ 6 TextBO Algorithm
After $T$ optimization iterations, we report as the algorithm’s output the sequence of best-performing outcomes
$\{\Phi(\pi_{t}^{j_{t}^{*}})\}_{t=1}^{T},\qquad j_{t}^{\ast}:=\operatorname{argmax}_{j}s_{t}^{j},$
i.e., the best trajectory at each iteration, according to the evaluation scores.
Note that, in each iteration, TextBO generates $N\times G$ prompt/outcome candidates for each trajectory before conducting one evaluation. This deliberate trade-off is the key to its evaluation efficiency: the algorithm uses inexpensive generation and critic-based comparisons to identify promising directions, so that each actual evaluation is performed on a small set of carefully selected, high-potential candidates.
7 Case Study: Digital Ad Optimization
We now instantiate our framework in a digital advertising setting, showing how an advertiser (or ad platform acting on behalf of an advertiser) can use TextBO to automate the ad generation–analysis–evaluation loop. Here, the goal is to refine the ad-generation prompts so that the resulting creatives are better aligned with a target population’s preference distribution, under a limited budget of costly ad evaluations. We specialize the general setup from $\S$3 to ad optimization as follows:
  • $\bullet$
    $\Phi$ is an ad-generation AI system that can generate ads given a prompt $\pi$. [^6] In practice, this image-generation model can be fine-tuned for a particular brand and/or customized to the platform. Our approach takes any pre-training as given.
  • $\bullet$
    For a prompt $\pi$, $\Phi(\pi)$ is the resulting ad creative deployed on the platform.
  • $\bullet$
    $\mathcal{X}$ is the population of platform users, and $\mathcal{D}_{\mathcal{X}}(\pi)$ is the (unknown) distribution of users who are exposed to $\Phi(\pi)$ when this ad is run.
  • $\bullet$
    $\mathcal{Y}$ is the space of observed ad responses (clicks, conversions, spend). For each user $x\in\mathcal{X}$, $y(\Phi(\pi),x)$ is the random ad response when $x$ is shown ad $\Phi(\pi)$.
  • $\bullet$
    $r:\mathcal{Y}\to[0,1]$ is a scoring rule that maps responses to a normalized effectiveness score (e.g., a weighted function of click, conversion, and spend).
The campaign objective is the same as the general objective introduced in $\S$3: $J(\pi)=\mathbb{E}\big[\,r\big(y(\Phi(\pi),x)\big)\big]$, where the expectation taken over $x\sim\mathcal{D}_{\mathcal{X}}(\pi)$ and the stochasticity in $y(\Phi(\pi),x)$.
7.1 Ad Campaign Scenarios
@@
We consider eight synthetic ad campaign scenarios that cover a diverse range of products across distinct categories, each defined by a creative brief that outlines the strategic and creative direction (see Web Appendix C.1 for the full creative briefs):
@@
  • Scenario 6: “MindGarden,” a subscription-based meditation and mindfulness app.
@@
7.2 Implementation Details
There are five main components of the algorithm: (1) initial prompt ($\pi_{0}$), and (2) Ad-generation module ($\Phi$), (3) Critic LLM and its four operators defined in $\S$5.1, (4) the population distribution $\mathcal{D}_{\mathcal{X}}(\pi)$ and the evaluation operator $\mathrm{Eval}()$, as defined in Equation (23). We now describe how each of these components is implemented in our experiments.
Initial prompts ($\pi_{0}^{j}$).
For each scenario, we first created 64 high-quality prompts from its creative brief using a meta-prompt. In Web Appendix C.2, see Figure A1 for the details of the meta-prompt and Figure A2 for an example of two prompts in scenario 1. Among those 64 ads, we chose five as initial prompts; we will describe the details of the choice rule in $\S$7.3.
Ad-generation module ($\Phi$).
We utilize a state-of-the-art image generation model, Google Imagen 4, [^7] See Vertex AI’s description (https://console.cloud.google.com/vertex-ai/publishers/google/model-garden/imagen-4.0-generate-preview-06-06) for details. Imagen 4 delivers high-fidelity, photorealistic text-to-image generation at up to 2K resolution, and excels in rendering intricate details such as realistic textures, lighting, and camera effects. Its features include typography accuracy for legible in-image text and precise adherence to complex prompts.
Critic LLM and its Operators.
Unless otherwise noted, we use Gemini 2.5 Flash  [^8] See Vertex AI’s description (https://docs.cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-5-flash) for details. as the multimodal critic model $M_{\text{critic}}$ and implement the four operators discussed in $\S$5.1 using the meta-prompts described in Web Appendix C.3: (1) PairwiseJudge operation; see meta-prompt in Figure A3, (2) MetaReflect operation; see meta-prompt in Figure A4, (3) Textual gradient operation $\nabla_{\text{text}}$; see meta-prompt in Figure A5, and (4) Rewrite operation Apply; see meta-prompt in Figure A6.
Evaluation module using persona data and LLM-based simulation.
We now describe how we implement the evaluation operator $\mathrm{Eval}$ in this case study. To evaluate how well an ad performs on the target population and to update prompts, we need (i) a target population ($\mathcal{D}_{\mathcal{X}}(\pi)$.) and (ii) a way to measure that population’s preferences for a given ad ($\mathrm{Eval}(\cdot)$). In principle, there are three ways define the evaluation operator – (1) field experiments/deployment in a real ad platform, (2) consumer surveys, and (3) simulated responses from a pre-defined target population. The former two options (field deployment and consumer surveys) are costly in time and resources, and can require large samples to identify effective differences across ads given low CTRs in digital ad settings; see (lewis2015unfavorable; johnson2023inferno). Thus, using these approaches for algorithm evaluation is not ideal.
@@ 7.2 Implementation Details
We split the Twin-2k-500 personas into 80% (1,647 personas) for training and 20% (411 personas) for testing. During algorithm training, each evaluation of an ad $\Phi(\pi)$ proceeds as follows: we randomly sample 200 personas from the training set and simulate ad effectiveness for each persona via the LLM. When we test a final ad (e.g., after running TextBO), we simulate its effectiveness for all personas in the test set.
We note that important advantage of our approach: using this persona distribution with LLM-generated preferences enables fast and controlled comparisons of different algorithms without large-scale field experiments or expensive surveys. [^10] Naive field deployment cannot be used to test the relative performance of algorithms, since an algorithm’s measured advantage would then mix its own effect with distributional effects induced by the ad platform’s targeting systems. Indeed, our approach can serve as a template for algorithm comparisons in many other settings where researchers require evaluations that reflect the variance/heterogeneity of human preferences. Further, we note that while LLM-based preference simulation may not reflect true human preferences (li2025llm; peng2025mega), it still induces a well-defined preference distribution on which we can benchmark optimization methods and algorithms. Specifically, kang2025llmpersonassubstitutefield proves that swapping humans for personas is equivalent to changing the evaluation population (e.g., from the New York population to the Jakarta population), under two benchmark hygiene conditions: (i) methods observe only the aggregate outcome (called aggregate-only observation condition) and (ii) evaluation depends only on the submitted artifact and not on the algorithm’s identity or provenance (called algorithm-blind evaluation condition). Furthermore, they show that simply increasing the size of the persona dataset is sufficient to guarantee that persona simulation becomes as useful a benchmark as field experimentation. Nevertheless, we also provide evaluations of our approach that do not leverage persona datasets; see $\S$8.4 for details.
7.3 Application of TextBO
We now describe how the TextBO algorithm is instantiated for the ad-optimization problem. Here, we choose the hyperparameters as: iterations $T=10$; trajectories $J=5$; gradient steps $G=5$; and candidates per gradient step $N=5$.
Initialization.
As described in $\S$7.2, for each scenario, we first generate 64 high-quality prompts from its creative brief using a meta-prompt. Among the 64 ads for each scenario, we identify the best initial ad, Best-of-64 (which will be used as one of the baselines but is not directly used in our algorithm; see $\S$8.1 for details), and the worst five, Worst5-of-64, by running a best-arm identification bandit algorithm (russo2016simple). Specifically, we alternate between a best-arm identification objective and a worst 5-arm identification objective over 5,000 sequential random samples from the training persona set. These Worst5-of-64 ads serve as starting points $\{\pi_{0}^{j}\}_{j=1}^{5}$ for TextBO procedure. This starting point allows us to attribute observed gains to the algorithms rather than to favorable starting prompts. [^11] To maintain a rigorous baseline, we make these 64 prompts detailed enough to yield strong creatives. This ensures that subsequent improvements are not merely due to trivial gains from starting with an under-specified prompts; instead, improvements reflect better alignment with the underlying target audience distribution. Let $\{s_{0}^{j}\}_{j=1}^{J}$ denote the evaluation scores for these five ads. Then, the tuples $\{(\pi_{0}^{j},\Phi(\pi_{0}^{j}),s_{0}^{j})\}_{j=1}^{J}$ form the initial history $H_{0}$ for the algorithm.
Best-of-N textual-gradient steps.
Within each optimization iteration $t$, TextBO refines each trajectory’s prompt $\pi_{t-1}^{j}$ through five steps of Best-of-N textual-gradients before evaluation (see Figure 3). We follow the generic procedure from $\S$6, but here specialize the critic’s context and edits to the ad domain.
[…]
  • $\bullet$
  • $\bullet$
  • $\bullet$
We then query these edits for the operation Apply operation to an initial prompt (see Figure A6 in Web Appendix C.3 for the meta-prompt) may yield candidates such as:
‘‘A photorealistic image of a sizzling plant-based burger with rich grill marks, served on a rustic wooden table with subtle greenery in the background and a small plant-based tag, conveying indulgence without sacrificing sustainability.’’
To select which candidate to keep for the next gradient step, TextBO applies the Best-of-N rule from $\S$5.2 to the ads $\{\Phi(\pi_{t,g}^{j,(i)})\}_{i=1}^{N}$. Specifically, following liu2025pairjudgermperformbestofn, we implement Best-of-N via a pairwise tournament: in each match, the critic compares two creatives side by side and predicts which will perform better using operator PairwiseJudge, given the campaign goals and the current reflection $R_{t-1}$ (see Figure A3 in Web Appendix C.3 for the meta-prompt). The tournament winner prompt and its corresponding ad, $\left(\pi_{t,g}^{j},\Phi\left(\pi_{t,g}^{j}\right)\right)$, are carried forward to the next gradient step. Repeating this for $g=1,\dots,G$ yields the final refined prompt $\pi_{t}^{j}$ and ad $\Phi(\pi_{t}^{j})$ for trajectory $j$ at iteration $t$.
Evaluation as the costly step.
After the $G$ gradient steps, TextBO performs a single expensive evaluation per trajectory (each of which consists of assessing the ad for 200 randomly sampled personas from the training persona set) by deploying $\Phi(\pi_{t}^{j})$ to get the evaluation result $s_{t}^{j}$ using the evaluation operator $\mathrm{Eval}$ from Equation (23) (instantiated using the evaluation module discussed in $\S$7.2). The new triple $(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})$ is appended to $H_{t}$, and the usual acceptance rule is applied: if $s_{t}^{j}\leq s_{t-1}^{j}$, the trajectory reverts to $(\pi_{t-1}^{j},\Phi(\pi_{t-1}^{j}),s_{t-1}^{j})$, ensuring non-decreasing scores along each trajectory.
Meta-reflection shared across trajectories.
After all $J$ trajectories have been evaluated at iteration $t$, the algorithm updates the global meta-reflection $R_{t}=\textsc{MetaReflect}(H_{t};M_{\text{critic}})$. Concretely, $M_{\text{critic}}$ is prompted to review a subset of the best- and worst-performing ads and prompts in $H_{t}$ and to summarize the key patterns that distinguish successful creatives from unsuccessful ones (see Figure A4 in Web Appendix C.3 for the meta-prompt). For example, it may infer that scenes with social eating and visible grill marks outperform isolated pack shots, or that emphasizing “satisfying” and “juicy” tends to improve effectiveness. This reflection is then injected into the context of subsequent textual-gradient generation and Best-of-N comparisons, allowing trajectories to share accumulated knowledge even as they explore different regions of the prompt space.
Reported outcome.
After $T=10$ iterations, the system reports the sequence of best-performing ads $\{\Phi\left(\pi_{t}^{j_{t}^{*}}\right)\}_{t=1}^{T}$, where $j_{t}^{\ast}=\arg\max_{j}s_{t}^{j}$ is the trajectory that achieved the highest score at iteration $t$.
8 Experiments for the Digital Ad optimization
In this section, we discuss the numerical performance of TextBO for the ad-optimization case (discussed in §7). In $\S$8.1, we first discuss the baseline algorithms we compare against, and then in $\S$8.2, we describe the experiment design. We present the main results in $\S$8.3 and then finally end with an ablation study in $\S$8.4.